OpenStudy (shelby1290):

Is (-1, 29) a solution of y= -2x^2 - 5x + 22? Show algebraically and explain

OpenStudy (trisaba):

\[\left( x, y \right)\] substitute

OpenStudy (shelby1290):

@Trisaba is this right? 29= -2(-1) ^2 -5(-1) + 22 29= -2-(-1) + 5 + 22 29= -1 + 5 + 22 29= 26 LS doesn't equal right side

OpenStudy (trisaba):

close they are not equal becaue 29 does not = 25

OpenStudy (shelby1290):

Okay thank you!